# The causal effects of modified treatment policies under network interference

### v2

@@ 1 Introduction
Scientists frequently seek to understand the causal effect of a given policy on an outcome $Y$ in a target population. Such policies commonly involve modifying the naturally occurring value of some continuous treatment or exposure $A$ for every population unit. Often, the scientist’s ideal goal is to answer the question, “how would the population mean of $Y$ change if the natural value of $A$ were increased or decreased?”
@@
Modified treatment policies (MTPs) (robins2004; munozPopulationInterventionCausal2012; haneuseEstimationEffectInterventions2013; young2014identification) are a class of interventions that define causal estimands well-suited for formulating such counterfactual questions about continuous exposures. An MTP can answer the question, “how much would the average value of $Y$ have changed had the natural value of $A$ been increased or decreased by an increment $\delta$?” where $\delta$ is chosen by the investigator. Hence, the MTP provides a causally interpretable generalization of the commonly applied procedure of estimating a regression coefficient. In fact, under specific structural assumptions, the effect of applying the MTP $A+1$ to every study unit is equivalent to the coefficient for the exposure $A$ in a main terms linear regression. Hence, though rarely mentioned explicitly by name, the causal effect of an MTP is a popular target estimand in studies that aim to evaluate policy effects on a population.
@@
The causal effect of an MTP carries several key advantages over alternative estimands. Unlike the causal dose-response curve, the MTP effect is identified even in settings where it may not be feasible—or, indeed, sensible—to deterministically set each unit’s exposure to the same level. Consequently, the MTP effect is identified even when certain exposure levels may suffer from non-overlap among sub-populations. Furthermore, the MTP effect estimand admits semi-parametric efficient estimators, accommodating the use of machine learning to flexibly capture nonlinear relationships and thus allowing the investigator to minimize the risks posed by model misspecification.
@@
Like other common causal inference methods, a key assumption of the classical MTP framework is non-interference—that one unit’s exposure does not impact any other units’ outcome (cox1958planning; rubin1980). This critical assumption cannot be made when, for example, study units correspond to geographic areas, such as counties or ZIP code areas. Here, the population units may move around; thus, assigning an exposure policy to a given county would affect not only those residing in that geographic region but also those commuting to and from said region. In general, ignoring interference between units invalidates identification and overtly introduces bias into analytic results (halloranDependentHappeningsRecent2016). Despite these challenges, data from studies that involve interference between study units can be exceedingly useful, even critical, to advancing science and policy in many settings of current interest. For example, environmental epidemiologists routinely leverage observational spatial data to study the effects of pollution on human health (elliottSpatialEpidemiologyCurrent2004; reichReviewSpatialCausal2021a; morrisonDefiningSpatialEpidemiology), a setting in which randomization is impractical and unethical, and where the “shifts in distribution” measured by MTPs are of interest (Tec2024).
@@
Contributions. We present several contributions to the literature on causal inference under network interference. Firstly, we introduce the concept of an induced MTP, a new type of intervention that identifies the causal effect of an MTP when network interference is present. This intervention accounts for how the application of an MTP to a given unit affects its neighbors in the known (or assumed) network structure. Secondly, applying the “coarea formula” from measure-theoretic calculus, we provide a novel identification result for the causal effect of an induced MTP in the network setting. Finally, we develop semi-parametric efficient estimators for the causal effect of an induced MTP by applying and building on the prior theoretical contributions of ogburnCausalInferenceSocial2022. Specifically, we resolve several of their methodological challenges, including (1) using cross-fitted machine learning for nuisance parameter estimation in place of restrictive parametric regression strategies; (2) deriving more tractable forms of the relevant nuisance parameters, which can be reliably estimated using standard regression tools; (3) eliminating reliance on computationally-intensive Monte Carlo procedures for estimation and inference; and (4) obtaining consistent variance estimators.
Outline. Section 2 reviews MTPs and causal inference under interference. Section 3 describes the induced MTP, including identification and estimation of causal effects. Section LABEL:section:sim-results reports numerical results verifying the proposed methodology, while Section LABEL:section:data-analysis discusses an illustrative data analysis in our motivating applied science context. We conclude and discuss future directions in Section LABEL:section:discussion.
@@ 2 Background
We will assume that the data-generating process can be expressed via a structural causal model (SCM, pearl2000models) encoding the temporal ordering between variables: $\mathbf{L}$ is generated first, then $\mathbf{A}$, and finally $\mathbf{Y}$. We denote by $Y(a)$ the counterfactual random variable (or potential outcome) generated by hypothetically intervening upon $A$ to set it to $a\in\mathcal{A}$ and allowing the impact of such an intervention to propagate downstream, to the component of the SCM that generates $Y$. Our goal is to reason about the causal relationship between $\mathbf{A}$ and $\mathbf{Y}$ in spite of the presence of confounders $\mathbf{L}$ and network interference between units $i=1,\ldots,n$.
@@ 2.1 Continuous Exposures with Modified Treatment Policies
A Modified Treatment Policy (MTP) is a user-specified function $d(a,l;\delta)$ that maps the observed value $a$ of an exposure $A$ to a new value and may itself depend on the natural (or pre-intervention) value of the exposure (haneuseEstimationEffectInterventions2013).
@@
MTPs can be used to define scientifically relevant causal estimands for continuous exposures. The population intervention causal effect of an MTP (munozPopulationInterventionCausal2012) is defined as $\mathbb{E}_{\mathsf{P}}[Y(d(A,L;\delta))-Y]$; that is, the average difference between the outcome $Y$ that did occur under the observed natural value of treatment $A$ and the counterfactual outcome $Y(d(A,L;\delta))$ that would have occurred under the investigator-supplied MTP $d(A,L;\delta)$. This estimand answers the scientific question, “what would happen if we applied, to the study population, a policy that modified the existing exposure according to a rule encoded by $d(\cdot;\delta)$?”
When $\delta=1$, the additive MTP (Example 1) carries the familiar interpretation often attributed to a linear regression coefficient; when $\delta=1.01$, the multiplicative MTP (Example 2) holds the same interpretation, but for a log-transformed exposure. Hence, MTPs may be seen as a non-parametric and causal extension of widely used associational estimands, formalizing the problem of quantifying how the mean counterfactual outcome would change under a shift in exposure value. MTPs allow the investigator to specify a wide range of interpretable interventions on continuous exposures that may be carried out in practice. The MTP framework has gained traction for its applicability in settings involving longitudinal, time-varying interventions (diazNonparametricCausalEffects2023; hoffman2024studying); causal mediation analysis (diaz2020causal; hejazi2023nonparametric), including with time-varying mediator–confounder feedback (gilbert2024identification); and causal survival analysis under competing risks (diaz2024causal). However, no work has, to our knowledge, extended the MTP framework to settings with dependent data characterized by interference between units.
@@ 2.2 Causal Inference Under Interference
Interference occurs when, for a given unit $i$, the outcome of interest $Y_{i}$ depends not only on its own assigned exposure $A_{i}$ but also upon the exposure $A_{j}$ of at least one other unit ($j\neq i$). Formally, we say interference occurs when $Y(a_{i},a_{j})\neq Y(a_{i},a_{j}^{\prime})$ if $a_{j}\neq a_{j}^{\prime}$. It is a component of the well-known stable unit treatment value assumption (SUTVA; rubin1980), commonly assumed for identification of causal estimands, including those of MTPs (haneuseEstimationEffectInterventions2013; young2014identification).
@@
Previous work has focused on settings exhibiting partial interference (hudgensCausalInferenceInterference2008; tchetgenCausalInferencePresence2012; halloranDependentHappeningsRecent2016), which occurs when units can be partitioned into clusters such that interference only occurs between units in the same cluster. Our work focuses instead on a broader setting, that of network interference (vanderlaanCausalInferencePopulation2014), which occurs when a unit’s outcome is subject to interference by other units’ exposures according to some arbitrary known network of relationships between units. When such interference is present, the data $\mathbf{O}$ includes an adjacency matrix or network profile, $\mathbf{F}$, describing each unit’s neighboring units, occasionally termed “friends” (sofryginSemiParametricEstimationInference2017).
@@
aronowEstimatingAverageCausal2017 demonstrate how to identify a causal estimand when SUTVA is violated due to network interference. To do this, they use an exposure mapping: a function that maps the exposure assignment vector $\mathbf{A}$ to the exposure actually received by each unit. The exposure received is a function of a unit’s original exposure $\mathbf{A}$ and covariates $\mathbf{L}$, including the network profile $\mathbf{F}$. If the exposure mapping is correctly specified and consistent, then SUTVA is restored, and causal effects subject to interference can be identified for the exposure arising under the exposure mapping.
@@
ogburnCausalInferenceSocial2022 and vanderlaanCausalInferencePopulation2014 rely on similar logic to identify population causal effects from data exhibiting a causally dependent structure, doing so by constructing “summary functions” of neighboring units’ exposures. Notably, ogburnCausalInferenceSocial2022 and vanderlaanCausalInferencePopulation2014 describe semi-parametric theory for efficiently estimating the effects of stochastic interventions in the network dependence setting. Stochastic interventions, which differ from MTPs, replace the natural value of exposure with a random draw from an investigator-supplied counterfactual distribution (munozPopulationInterventionCausal2012). While a mathematically elegant strategy, the interpretation of a stochastic intervention is typically challenging—at times, even impractical—as real-world policies can seldom be defined by randomly assigning (post-intervention) exposure values to study units. Furthermore, the estimation of their causal effects is challenging: previous works in the dependent data setting have been restricted to relatively bespoke and restrictive parametric modeling of nuisance parameters, coupled with Monte Carlo procedures for point and variance estimation.
@@
Notably, while the hypothetical exposure that results from this random draw is not guaranteed to match that which would result from an MTP, the two classes of interventions may be constructed to yield equivalent counterfactual means (young2014identification). Given the similarities between these intervention schemes, we extend recent theoretical developments to construct semi-parametric efficient estimators of the effects of MTPs under network interference. Our work reveals that, for MTPs, much of the previously established semi-parametric theory can be reduced in ways that simplify the application of machine learning or non-parametric regression for nuisance estimation.
@@
But what does it mean to intervene on a summary function? If one were to intervene on the summary directly, the resulting collection of counterfactual exposures could plausibly be inconsistent with the structure of the network. Some existing works seek to circumvent this issue by recasting the desired estimand as a mean of individual-level causal effects (aronowEstimatingAverageCausal2017; athey2017; Savje2023). When investigators aim to estimate the impact of a hypothetical policy, however, this strategy will not answer questions of scientific interest—for the estimand does not correspond to a population-level intervention that could be implemented in practice. In order to estimate a population-level causal effect, one must consider first intervening, and only then applying the summary or exposure mapping—a process more naturally applicable to MTPs.
@@
Other relevant works address interference under different assumptions: random networks (clarkCausalInferenceStochastic2024), multiple outcomes (shin2023), long-range dependence (tchetgentchetgenAutoGComputationCausalEffects2021; tchetgen2025autodrestimation), bipartite graphs (ziglerBipartiteInterferenceAir2023), and unknown network structure (ohnishi2022degree; hoshino2023causal). We build on the setting described by ogburnCausalInferenceSocial2022, as their scientific goals most closely resemble those of the MTP framework.
@@ 3 Methodology
Suppose there exists a network describing whether two units are causally dependent with adjacency matrix $\mathbf{F}$, where $F_{i}$ denotes the neighbors of unit $i$. For each unit $i$, a set of confounders $L_{i}$ is drawn, followed by an exposure $A_{i}$ based on a summary $L_{i}^{s}$ of its own and its neighbors’ confounders, and finally an outcome based on $L_{i}^{s}$ and a summary of its own and its neighbors’ exposures, $A_{i}^{s}$. This data-generating process can be defined formally as the SCM in Equation (4):
@@
Following ogburnCausalInferenceSocial2022, we assume error vectors $(\varepsilon_{L_{1}},\ldots,\varepsilon_{L_{n}})$, $(\varepsilon_{A_{1}},\ldots,\varepsilon_{A_{n}})$, and $(\varepsilon_{Y_{1}},\ldots,\varepsilon_{Y_{n}})$ are independent of each other, with entries identically distributed and either $\varepsilon_{i}\mbox{$\perp\!\!\!\perp$}\varepsilon_{j}$ provided $\{i,j\}\not\subseteq F_{k},\,\forall\,\,k\in 1,\ldots,n$ or $\text{Cov}(\varepsilon_{i},\varepsilon_{j})\geq 0$ otherwise. That is, errors between units are independent provided that the units are neither directly connected nor share ties with a common node in the interference network given by $\mathbf{F}$; otherwise, errors across units may be positively correlated. Positive correlation ensures Theorem LABEL:theorem:ogburn will hold, and is typical in applied settings with interference.
Interference bias arises when the data arise from the SCM (4) but investigators wrongly assume that $f_{Y}$ is a function only of $A_{i}$ and $L_{i}$, and not of $\{A_{j}\colon F_{ij}\neq 0\}$ or $\{L_{j}\colon F_{ij}\neq 0\}$. Since interference violates the consistency rule (Pearl2010), commonly relied upon to identify causal effects, ignoring its presence, even inadvertently, leads to a failure in identification and consequently risks biased estimation. Under the SCM (4), identifiability of the causal effect of applying an exposure to all units $A_{j}\colon j\in F_{i}$ can be restored by controlling for all $L_{j}\colon F_{ij}\neq 0$ directly or via the dimension-reducing summaries $A_{i}^{s}$ and $L_{i}^{s}$ (vanderlaanCausalInferencePopulation2014) of a unit’s neighbors’ confounders and exposures.
Since $A_{i}^{s}$ and $L_{i}^{s}$ are not directly observed but are functions that follow from the scientific problem at hand, we denote (in a slight abuse of notation) $A_{i}^{s}=s_{F_{i}}(\mathbf{A},\mathbf{L})$ and $L_{i}^{s}=s_{F_{i}}(\mathbf{L})$, which we use to denote that, for unit $i$, the summary function depends on $i$’s column of the network adjacency matrix $\mathbf{F}$. This means that $A_{i}^{s}$ and $L_{i}^{s}$ depend only on neighboring units. As just one example, $A_{i}^{s}$ may be a (possibly weighted) sum of neighbors: $s_{F_{i}}(\mathbf{A},\mathbf{L})=\sum_{j=1}^{n}F_{ij}A_{j}=\sum_{F_{ij}\neq 0}A_{j}$.

### v3

@@ 1 Introduction
Scientists frequently seek to understand the causal effect of a given policy on an outcome $Y$ in a target population. Such policies commonly involve modifying the natural value of some continuous treatment or exposure $A$; that is, the value $A$ would have taken had its distribution been left unaltered (Sarvet & Stensrud, 2025). Often, the scientist’s ideal goal is to answer the question, “how would the population mean of $Y$ change if the natural value of $A$ were increased or decreased?”
@@
Modified treatment policies (MTPs) (Robins et al., 2004; Díaz & van der Laan, 2012; Haneuse & Rotnitzky, 2013; Young et al., 2014) are a class of interventions that define causal estimands well-suited for formulating such counterfactual questions about continuous exposures. An MTP can answer the question, “how much would the average value of $Y$ have changed had the natural value of $A$ been increased or decreased by an increment $\delta$?” where $\delta$ is chosen by the investigator. Hence, the MTP provides a causally interpretable generalization of the commonly applied procedure of estimating a regression coefficient. In fact, under specific structural assumptions, the effect of applying the MTP $A+1$ to every study unit is equivalent to the coefficient for the exposure $A$ in a main terms linear regression. Hence, though rarely mentioned explicitly by name, the causal effect of an MTP is a popular target estimand in studies that aim to evaluate policy effects on a well-defined population.
@@
The causal effect of an MTP carries several key advantages over alternative estimands. Unlike the causal dose-response curve, the MTP effect is identified even in settings where it may not be feasible—or, indeed, sensible—to deterministically set each unit’s exposure to the same level. Consequently, the MTP effect is identified even when certain exposure levels may suffer from non-overlap among sub-populations. Furthermore, the MTP effect estimand admits standard semi-parametric efficient estimators, facilitating the use of machine learning to flexibly capture nonlinear relationships and thus allowing the investigator to mitigate the risks posed by model misspecification.
@@
Like other common causal inference methods, a key assumption of the classical MTP framework is non-interference—that one unit’s exposure does not impact any other units’ outcome (Cox, 1958; Rubin, 1980). This critical assumption cannot be made when, for example, study units correspond to geographic areas, such as counties or ZIP code areas. Here, the population units may move around; thus, assigning an exposure policy to a given county would affect not only those residing in that geographic region but also those commuting to and from said region. In general, ignoring interference between units invalidates identification and overtly introduces bias into analytic results (Halloran & Hudgens, 2016). Despite these challenges, data from studies that involve interference between study units can be exceedingly useful, even critical, to advancing science and policy in many settings of current interest. For example, environmental epidemiologists routinely leverage observational spatial data to study the effects of pollution on human health (Elliott & Wartenberg, 2004; Reich et al., 2021; Morrison et al., 2024), a setting in which randomization is impractical and unethical, and where the “shifts in distribution” measured by MTPs are of interest (Tec et al., 2024).
@@
Contributions. We present several contributions to the literature on causal inference under network interference. Firstly, we introduce the concept of an induced MTP, a new type of intervention that identifies the causal effect of an MTP when network interference is present. This intervention accounts for how the application of an MTP to a given unit affects its neighbors in the known (or assumed) network structure. Secondly, applying the “coarea formula” from measure-theoretic calculus, we provide a novel identification result for the causal effect of an induced MTP in the network setting. Finally, we develop semi-parametric efficient estimators for the causal effect of an induced MTP by applying and building on the prior theoretical contributions of Ogburn et al. (2022). Specifically, we resolve several of their methodological challenges, including (1) using cross-fitted machine learning for nuisance parameter estimation in place of restrictive parametric regression strategies; (2) deriving more tractable forms of the relevant nuisance parameters, which can be reliably estimated using standard regression tools; (3) eliminating reliance on computationally-intensive Monte Carlo procedures for estimation and inference; and (4) obtaining consistent variance estimators.
Outline. Section 2 reviews MTPs and causal inference under interference. Section 3 describes the induced MTP, including identification and estimation of causal effects. Section 4 reports numerical results verifying the proposed methodology, while Section 5 discusses an illustrative data analysis in our motivating applied science context. We conclude and discuss future directions in Section 6.
@@ 2 Background
We will assume that the data-generating process can be expressed via a structural causal model (SCM, Pearl (2000)) encoding the temporal ordering between variables: $\mathbf{L}$ is generated first, then $\mathbf{A}$, and finally $\mathbf{Y}$. We denote by $Y(a)$ the counterfactual random variable (or potential outcome) generated by hypothetically intervening upon $A$ to set it to $a\in\mathcal{A}$ and allowing the impact of such an intervention to propagate downstream, to the component of the SCM that generates $Y$. Our goal is to reason about the causal relationship between $\mathbf{A}$ and $\mathbf{Y}$ in spite of the presence of confounders $\mathbf{L}$ and network interference between units $i=1,\ldots,n$.
@@ 2.1 Continuous Exposures with Modified Treatment Policies
A Modified Treatment Policy (MTP) is a user-specified function $d(a,l;\delta)$ that maps the observed value $a$ of an exposure $A$ to a new value and may itself depend on the natural (or pre-intervention) value of the exposure (Haneuse & Rotnitzky, 2013).
@@
MTPs can be used to define scientifically relevant causal estimands for continuous exposures. The population intervention causal effect of an MTP (Díaz & van der Laan, 2012) is defined as $\mathbb{E}_{\mathsf{P}}[Y(d(A,L;\delta))-Y]$; that is, the average difference between the outcome $Y$ that did occur under the observed natural value of treatment $A$ and the counterfactual outcome $Y(d(A,L;\delta))$ that would have occurred under the investigator-supplied MTP $d(A,L;\delta)$. This estimand answers the scientific question, “what would happen if we applied, to the study population, a policy that modified the existing exposure according to a rule encoded by $d(\cdot;\delta)$?”
When $\delta=1$, the additive MTP (Example 1) carries the familiar interpretation often attributed to a linear regression coefficient; when $\delta=1.01$, the multiplicative MTP (Example 2) holds the same interpretation, but for a log-transformed exposure. Hence, MTPs may be seen as a non-parametric and causally interpretable extension of widely used associational estimands, formalizing the problem of quantifying how the mean counterfactual outcome would change under a shift in exposure value. MTPs allow the investigator to specify a wide range of interpretable interventions on continuous exposures that may be carried out in practice. The MTP framework has gained traction for its applicability in settings involving longitudinal, time-varying interventions (Díaz et al., 2023; Hoffman et al., 2024); causal mediation analysis (Díaz & Hejazi, 2020; Hejazi et al., 2023), including with time-varying mediator–confounder feedback (Gilbert et al., 2024); and causal survival analysis under competing risks (Díaz et al., 2024). However, no work has, to our knowledge, extended the MTP framework to settings with dependent data characterized by interference between units.
@@ 2.2 Causal Inference Under Interference
Interference occurs when, for a given unit $i$, the outcome of interest $Y_{i}$ depends not only on its own assigned exposure $A_{i}$ but also upon the exposure $A_{j}$ of at least one other unit ($j\neq i$). Formally, we say interference occurs when $Y(a_{i},a_{j})\neq Y(a_{i},a_{j}^{\prime})$ if $a_{j}\neq a_{j}^{\prime}$. It is a component of the well-known stable unit treatment value assumption (SUTVA; Rubin, 1980), commonly assumed for identification of causal estimands, including those of MTPs (Haneuse & Rotnitzky, 2013; Young et al., 2014).
@@
Previous work has focused on settings exhibiting partial interference (Hudgens & Halloran, 2008; Tchetgen Tchetgen & VanderWeele, 2012; Halloran & Hudgens, 2016), which occurs when units can be partitioned into clusters such that interference only occurs between units in the same cluster. Our work focuses instead on a broader setting, that of network interference (van der Laan, 2014), which occurs when a unit’s outcome is subject to interference by other units’ exposures according to some arbitrary known network of relationships between units. When such interference is present, the data $\mathbf{O}$ includes an adjacency matrix or network profile, $\mathbf{F}$, describing each unit’s neighboring units, occasionally termed “friends” (Sofrygin & van der Laan, 2017).
@@
Aronow & Samii (2017) demonstrate how to identify a causal estimand when SUTVA is violated due to network interference. To do this, they use an exposure mapping: a function that maps the exposure assignment vector $\mathbf{A}$ to the exposure actually received by each unit. The exposure received is a function of a unit’s original exposure $\mathbf{A}$ and covariates $\mathbf{L}$, including the network profile $\mathbf{F}$. If the exposure mapping is correctly specified and consistent, then SUTVA is restored, and causal effects subject to interference can be identified for the exposure arising under the exposure mapping.
@@
Ogburn et al. (2022) and van der Laan (2014) rely on similar logic to identify population causal effects from data exhibiting a causally dependent structure, doing so by constructing “summary functions” of neighboring units’ exposures. Notably, they describe semi-parametric theory for efficiently estimating the effects of stochastic interventions in the network dependence setting. Stochastic interventions, which differ from MTPs, replace the natural value of exposure with a random draw from an investigator-supplied counterfactual distribution (Díaz & van der Laan, 2012). While a mathematically elegant strategy, the interpretation of a stochastic intervention is typically challenging—at times, even impractical—as real-world policies can seldom be defined by randomly assigning (post-intervention) exposure values to study units. Furthermore, this setting gives rise to estimation challenges: empirical evaluations in Ogburn et al. (2022) and related work by Sofrygin & van der Laan (2017) relied on relatively bespoke parametric modeling of nuisance parameters, coupled with Monte Carlo procedures for point and variance estimation.
@@
Notably, while the hypothetical exposure that results from this random draw is not guaranteed to match that which would result from an MTP, the two classes of interventions may be constructed to yield equivalent counterfactual means (Young et al., 2014). Given the similarities between these intervention schemes, we extend recent theoretical developments to construct semi-parametric efficient estimators of the effects of MTPs under network interference. Our work reveals that, for MTPs, much of the previously established semi-parametric theory can be reduced in ways that simplify the application of machine learning and adaptive, non-parametric regression for nuisance estimation.
@@
But what does it mean to intervene on a summary function? If one were to intervene on the summary directly, the resulting collection of counterfactual exposures could plausibly be inconsistent with the structure of the network. Some existing works seek to circumvent this issue by recasting the desired estimand as a mean of individual-level causal effects (Aronow & Samii, 2017; Athey et al., 2017; Sävje, 2023). When investigators aim to estimate the impact of a hypothetical policy, however, this strategy will not answer questions of scientific interest—for the estimand does not correspond to a population-level intervention that could be implemented in practice. In order to estimate a population-level causal effect, one must consider first intervening, and only then applying the summary or exposure mapping—a process more naturally applicable to MTPs.
@@
Other relevant works address interference under different assumptions: random networks (Clark & Handcock, 2024), multiple outcomes (Shin et al., 2023), long-range dependence (Tchetgen Tchetgen et al., 2021; Liu et al., 2025), bipartite graphs (Zigler et al., 2023), and unknown network structure (Ohnishi et al., 2022; Hoshino & Yanagi, 2023). We build on the setting described by Ogburn et al. (2022), as their scientific goals most closely resemble those of the MTP framework.
@@ 3 Methodology
Suppose there exists a network describing whether two units are causally dependent with adjacency matrix $\mathbf{F}$, where $F_{i}$ denotes the neighbors of unit $i$. For each unit $i$, a set of confounders $L_{i}$ is drawn, followed by an exposure $A_{i}$ based on a summary $L_{i}^{s}$ of its own and its neighbors’ confounders, and finally an outcome $Y_{i}$ based on $L_{i}^{s}$ and a summary of its own and its neighbors’ exposures, $A_{i}^{s}$. This data-generating process can be defined formally as the SCM in Equation (4):
@@
Since $A_{i}^{s}$ and $L_{i}^{s}$ are not directly observed but are functions that follow from the scientific problem at hand, it follows that (in a slight abuse of notation) $A_{i}^{s}=s_{F_{i}}(\mathbf{A},\mathbf{L})$ and $L_{i}^{s}=s_{F_{i}}(\mathbf{L})$, which we use to denote that, for unit $i$, the summary function depends on $i$’s column of the network adjacency matrix $\mathbf{F}$. This means that $A_{i}^{s}$ and $L_{i}^{s}$ depend only on neighboring units. As just one example, $A_{i}^{s}$ may be a (possibly weighted) sum of neighbors: $s_{F_{i}}(\mathbf{A},\mathbf{L})=\sum_{j=1}^{n}F_{ij}A_{j}=\sum_{F_{ij}\neq 0}A_{j}$.
Following Ogburn et al. (2022), we assume error vectors $(\varepsilon_{L_{1}},\ldots,\varepsilon_{L_{n}})$, $(\varepsilon_{A_{1}},\ldots,\varepsilon_{A_{n}})$, and $(\varepsilon_{Y_{1}},\ldots,\varepsilon_{Y_{n}})$ are independent of each other, with entries identically distributed and either $\varepsilon_{i}\mbox{$\perp\!\!\!\perp$}\varepsilon_{j}$ provided $\{i,j\}\not\subseteq F_{k},\,\forall\,\,k\in 1,\ldots,n$ or $\text{Cov}(\varepsilon_{i},\varepsilon_{j})\geq 0$ otherwise. That is, errors between units are independent provided that the units are neither directly connected nor share ties with a common node in the interference network given by $\mathbf{F}$; otherwise, errors across units may be positively correlated. Positive correlation ensures Theorem 1 will hold, and is typical in applied settings with interference.
Interference bias arises when the data arise from the SCM (4) but investigators wrongly assume that $f_{Y}$ is a function only of $A_{i}$ and $L_{i}$, and not of $\{A_{j}\colon F_{ij}\neq 0\}$ or $\{L_{j}\colon F_{ij}\neq 0\}$. Since interference violates the consistency rule (Pearl, 2010), commonly relied upon to identify causal effects, ignoring its presence, even inadvertently, leads to a failure in identification and consequently risks biased estimation. Under the SCM (4), identifiability of the causal effect of applying an exposure to all units $A_{j}\colon j\in F_{i}$ can be restored by controlling for all $L_{j}\colon F_{ij}\neq 0$ directly or via the dimension-reducing summaries $A_{i}^{s}$ and $L_{i}^{s}$ (van der Laan, 2014) of a unit’s neighbors’ confounders and exposures.